# pnasresearcharticle

### v1

@@ Results
Our results so far have dealt with strong reciprocal CS, $k_{\rm CS}=0.05$. If the strength of CS diminishes ($k_{\rm CS}$ grows), the extinction probabilities at the end of the treatment decrease, as expected given the previous discussion (Fig. 4A). Although the dependence of extinction probability on $\tau$ is qualitatively similar, i.e. optimal $\tau$ remain the same throughout, the maximum extinction probability decreases monotonically as $k_{\rm CS}$ increases, and for $k_{\rm CS}>0.3$ it is approximately zero for all $\tau$. That is, for subinhibitory concentrations, CS is a necessary condition for the success of sequential therapies.
@@
We wondered then whether this dependency on CS would be weakened if we increased antibiotic doses. We reasoned that, as $k_{A},k_{B}\to 0$, the total extinction probability should increase. For concentrations close to the MIC, i.e. $k=0.1$ or lower, a threshold behavior emerges where populations go extinct regardless of the switching period (Fig. 4B,C), suggesting that extinction is driven primarily by the strength of inhibition rather than the switching dynamics. Indeed, this behavior persists even in the absence of CS (SI Appendix, Fig. S6). However, for lower antibiotic concentrations $k>0.1$ we observe a unimodal dependency of the extinction probability on the dose: for a fixed $\tau$, extinction probabilities go up as we decrease $k$ from $1$, and then decrease after a certain dose (Fig. 4B,C). We reason that this is due to a change in population dynamics: when antibiotic inhibition increases, the single-resistant mutant quickly dominates the population, which then becomes extinct more easily as we switch the antibiotic. However, if antibiotic inhibition crosses a given threshold, the wild type population is so low that the influx of resistant mutants actually becomes slower, and therefore the population does not become extinct after switching. We support this qualitative reasoning looking at the mean population before the first switch (SI Appendix, Fig. S7). In the same way, extinction probabilities show a unimodal dependency on mutation: as $\mu_{1}$ increases, extinction probabilities go up, again due to an increase in the abundance of single-resistant mutants (SI Appendix, Fig. S8). However, after a critical threshold, the extinction probability starts to decrease, in this case because the double-resistant strain emerges in the population, while the susceptible strain concurrently exhibits a recovery (SI Appendix, Fig. S9). Moreover, as $\mu_{2}/\mu_{1}$ increases, the extinction probability decreases more markedly, which is consistent with the accelerated recovery of the susceptible strain under these conditions, since single-resistant mutants become less frequent when $\mu_{2}$ is relatively larger.
Figure 5: Extinction probability curves for different final simulation times. (A) $t_{\text{end}}=150$, $t_{\text{end treatment}}=100$. (B) $t_{\text{end}}=250$, $t_{\text{end treatment}}=200$. (C) $t_{\text{end}}=350$, $t_{\text{end treatment}}=300$. (D) $t_{\text{end}}=450$, $t_{\text{end treatment}}=400$.
Optimal switching periods
After providing an evolutionary, population-dynamics argument for the success of sequential therapies based on strong reciprocal CS, and studying the effect of varying several parameters, we return to finding the optimal switching periods. In mathematical terms, and following our previous discussion, the function we want to maximize is $E(\tau)=1-\big[1-p(\tau)\big]^{n}$ where $p(\tau)$ is an increasing function of $\tau$ (Fig. 2E) and the number of switches $n=\lfloor T/\tau\rfloor$ is also a function of $\tau$ for a given treatment duration $T$. For a fixed $T$, there are two opposing trends affecting $E(\tau)$, as longer $\tau$ increases $p(\tau)$, but decreases $n$. Note also that, independent of $T$, the optimal therapy for a given number of switches always appears at $\tau=T/n$, i.e. when the time before switching is longest. In other words, we only need to discuss optimal therapies in terms of number of switches and treatment duration. For instance, for $T=100$ (Fig. 5A) the optimal therapy is either one or two switches, corresponding to $\tau=100$ and $\tau=50$ respectively. However, a closer analysis of longer therapies reveals that intermediate $\tau$ outperform longer ones as $p(\tau)$ saturates but the number of switches is higher: for $T=200$ (Fig. 5B) the optimal therapy is either two ($\tau=100$) or three ($\tau\approx 66.7$) switches, whereas for $T=300$ the maximum is found at three ($\tau=100$) or four ($\tau=75$) switches (Fig. 5C). For $T=400$, the interval between $\tau=50$ and $\tau=100$ yields comparable extinction probabilities (Fig. 5D). We could use our heuristic for $p(\tau)$ to obtain the optimal $\tau$ for a given $T$. However, here we are concerned with giving a qualitative understanding of optimal switching periods, as the actual value will depend on population parameters such as mutation and death rates and should be estimated for actual pathogens. Crucially, the optimal $\tau$ depends nontrivially on treatment duration $T$, a fact that has to be taken into account when designing a therapy.
Discussion
@@
The use of collateral sensitivity (CS) to design sequential therapies has received widespread attention in recent years (17, 34). The underlying rationale is that antibiotics can be used to steer populations toward genotypic states exhibiting CS, thereby increasing the efficacy of subsequent treatments (35). In this work, we develop a simple mathematical framework that captures key features of sequential therapies, enabling quantitative exploration of extinction dynamics and providing support for prior evolutionary hypotheses. While experimental studies have demonstrated the potential of sequential therapies to suppress resistance (20, 36, 21), theoretical efforts have largely relied on deterministic models where extinctions do not occur (37, 38, 39), although a recent work has used stochastic modeling to explore the effect of antibiotic pulses (24). In contrast to deterministic models, our stochastic modeling framework captures extinction dynamics directly. This enables the identification of optimal treatment strategies based on true eradication events and allows us to explore how extinction probabilities depend on switching timing, antibiotic potency, and mutation rates.
@@
In (37), Beardmore and Peña-Miller state that a successful switching strategy, based on clinical observations (40), is “if the observed level of resistance to an antibiotic is too high, exchange it for a different antibiotic”, which is fully consistent with our results. In fact, a key result from our analysis is that fast sequential therapies are suboptimal, and that we need to allow for the evolution of resistance above a threshold in order to eradicate the population after the switch. The tension between waiting long enough for the single-resistant mutant to dominate the population and maximizing the number of switches leads to different optimal switching periods depending on the treatment duration (Fig. 5).
@@
Furthermore, we show that, under our theoretical framework, we need strong CS for our therapies to work. While this seems a rather stringent condition, the good news is that such a therapy will work even under subinhibitory concentrations, which will always appear within the human body as a result of diffusion through tissues (41).
@@
In addition, we find a unimodal dose-extinction relationship (Fig. 4C), which has been observed before in both experiments (33) and models (42) and which can be fully explained using population dynamics arguments: we need single-resistant mutants to dominate the population; these are selected as antibiotic concentration increases, but selection can be hampered if doses are too high. We observe a similar relationship with mutation rates (SI Appendix, Fig. S8), also supported by population dynamics arguments. This result suggests that increasing mutation rates might be a good complement to this kind of therapies, up to a point where we start facilitating the evolution of double-resistants. This suggestive strategy should be explored with care and checked experimentally before making any therapy recommendations.
Several limitations of our work warrant discussion. First, our model considers only four genotypes and two antibiotics with symmetric effects; real bacterial populations exhibit more complex collateral sensitivity networks (17, 18, 19) and heterogeneous pharmacokinetics/pharmacodynamics (39). Second, CS patterns need not (and in fact do not) remain constant through treatment (43) and this will pose a serious hindrance for translating our evolutionary insights into actual therapies. We should, therefore, test our hypotheses in the laboratory, and we have already identified potential candidates that we can test in vitro. Notably, the pair ciprofloxacin-tobramycin has been shown to robustly evolve CS mutants in P. aeruginosa, and double-resistance rarely evolves (19, 44, 45), so it should be a proper experimental setup to test our hypotheses.
A full treatment of optimal switching therapies including nonperiodic switches, as done in (37, 38), was out of the scope of the present paper, but we intend to explore it in more depth in the future: our analysis in this paper suggest that we could implement control‐theoretic optimization of switching schedules, potentially in real time using patient‐specific bacterial load measurements, an exciting direction. Ultimately, our results support the promise of sequential therapies guided by collateral sensitivity and mathematical modeling as a complementary strategy in the fight against antimicrobial resistance.
[…]
@@
[…]

### v3

@@ III.2 Heuristic explanation for the change in extinction probabilities
Our results so far have dealt with strong reciprocal CS, $k_{\rm CS}=0.05$. If the strength of CS diminishes ($k_{\rm CS}$ grows), the extinction probabilities at the end of the treatment decrease, as expected given the previous discussion (Fig. 5(a)). Although the dependence of extinction probability on $\tau$ is qualitatively similar, i.e. optimal $\tau$ remain the same throughout, the maximum extinction probability decreases monotonically as $k_{\rm CS}$ increases, and for $k_{\rm CS}>0.3$ it is approximately zero for all $\tau$. That is, for subinhibitory concentrations, CS is a necessary condition for the success of sequential therapies.
@@
We wondered then whether this dependency on CS would be weakened if we increased antibiotic doses. We reasoned that, as $k_{A},k_{B}\to 0$, the total extinction probability should increase. For concentrations close to the minimum inhibitory concentration (MIC), i.e. $k=0.1$ or lower, a threshold behavior emerges where populations go extinct regardless of the switching period (Fig. 5(b,c)), suggesting that extinction is driven primarily by the strength of inhibition rather than the switching dynamics. Indeed, this behavior persists even in the absence of CS (Appendix F, Fig. 12). However, for lower antibiotic concentrations $k>0.1$ we observe a unimodal dependency of the extinction probability on the dose: for a fixed $\tau$, extinction probabilities go up as we decrease $k$ from $1$, and then decrease after a certain dose (Fig. 5(b,c)). We reason that this is due to a change in population dynamics: when antibiotic inhibition increases, the single-resistant mutant quickly dominates the population, which then becomes extinct more easily as we switch the antibiotic. However, if antibiotic inhibition crosses a given threshold, the wild type population is so low that the influx of resistant mutants actually becomes slower, and therefore the population does not become extinct after switching. We support this qualitative reasoning looking at the mean population before the first switch (Appendix F, Fig. 13). In the same way, extinction probabilities show a unimodal dependency on mutation: as $\mu_{1}$ increases, extinction probabilities go up, again due to an increase in the abundance of single-resistant mutants (Appendix F, Fig. 14). However, after a critical threshold, the extinction probability starts to decrease, in this case because the double-resistant strain emerges in the population, while the susceptible strain concurrently exhibits a recovery (Appendix F, Fig. 15). Moreover, as $\mu_{2}/\mu_{1}$ increases, the extinction probability decreases more markedly, which is consistent with the accelerated recovery of the susceptible strain under these conditions, since single-resistant mutants become less frequent when $\mu_{2}$ is relatively larger.
III.4 Optimal switching periods
After providing an evolutionary population-dynamics argument for the success of sequential therapies based on strong reciprocal CS, and studying the effect of varying several parameters, we return to finding the optimal switching periods.
Figure 6: Optimal therapy. (a) Extinction probability as a function of the switching period. (b) Probability of double resistant mutants at the end of the treatment, conditioned on the population not being extinct. (c) Pareto front of optimal switching therapies. The crosses indicate therapies that belong to the Pareto front, while the circles show dominated (suboptimal) therapies. The grey dashed line indicates the empirical fit. The probability that a therapy is in the Pareto front is maximized for $\tau\approx 42$. Simulated therapy trajectories were generated with different switching periods ($\tau$) and numbers of treatment rounds ($n$), constrained such that $\tau\cdot n+50<350$, number of trajectories =10,000=10,000, $k=0.5$, $k_{CS}=0.05$.
Figure 7: Optimal window. Simulated therapy trajectories were generated with different switching periods ($\tau$) and numbers of treatment rounds ($n$), constrained such that $\tau\cdot n+50<350$. (a) Fraction of therapies with a given switching period $\tau$ that lie on the Pareto front. (b) Distribution of these fractions across the different numbers of treatment rounds. We observe that certain switching periods consistently fall on the Pareto front regardless of the number of rounds $\tau\approx 42$, while others are optimal only for one $\tau\approx 20$ or three $\tau\approx 90$ switching events. Number of trajectories =10000=10000.
So far, we have discussed optimality only in terms of maximizing the cumulative extinction probability at the end of the treatment. In mathematical terms, this function is $E(\tau,n)=1-\big[1-p(\tau)\big]^{n}$ where $p(\tau)$ is an increasing function of $\tau$, as we have discussed (Fig. 2(e)). Therefore, $E(\tau,n)$ increases both with $\tau$ and with the number of switches $n$: if the population does not become extinct after the first change, it may do so after subsequent changes. In fact, for a fixed $\tau$, the probability of extinction can increase as much as we want by increasing $n$ (Fig. 6(a)).
However, this comes with a cost: longer therapies also have more probability that the double-resistant genotype $x_{3}$ appears in the population at the end of the treatment (Fig. 6(b)), so if the population is not extinct it will become resistant. Hence, there is a trade-off between maximizing extinction probabilities and minimizing the probability of double-resistance, leading to a Pareto front of feasible therapies (Fig. 6(c)) that optimize both objectives. For a fixed extinction probability $E(\tau,n)$, the Pareto front contains those therapies that minimize the probability of double resistance at the end of the treatment, conditioned on the population not being extinct, which we define as $R(\tau,n)=\Pr(N_{3}(T)\geq 1|N(T)>0)$. If a therapy is not on the Pareto front, we can always modify it so that either $E(\tau,n)$ or $1-R(\tau,n)$ are improved: they are dominated by other, better therapies. Crucially, therapies on the Pareto front cannot improve $E(\tau,n)$ without increasing $R(\tau,n)$ as well. We numerically find that the Pareto front is captured by the curve $E(\tau,n)^{1.09}+(1-R(\tau,n))^{1.22}=1$ (Fig. 6(c)). We can also show that the probability of being in the Pareto front is very high for $\tau\approx 42$ (Fig. 7(a)), independently of the number of changes (Fig. 7(b)) and lower everywhere else. Hence, for the parameter set studied in this paper, choosing $\tau$ around that value will lead to optimal therapies: we can then choose whether we want to maximize extinction or minimize resistance by changing the number of switches $n$.
IV Discussion
@@
The use of collateral sensitivity (CS) to design sequential therapies has received widespread attention in recent years [17, 36]. The underlying rationale is that antibiotics can be used to steer populations toward genotypic states exhibiting CS, thereby increasing the efficacy of subsequent treatments [37]. In this work, we develop a simple mathematical framework that captures key features of sequential therapies, enabling quantitative exploration of extinction dynamics and providing support for prior evolutionary hypotheses. While experimental studies have demonstrated the potential of sequential therapies to suppress resistance [20, 38, 21], theoretical efforts have largely relied on deterministic models where extinctions do not occur [39, 40, 41], although a recent work has used stochastic modeling to explore the effect of antibiotic pulses [24]. In contrast to deterministic models, our stochastic modeling framework captures extinction dynamics directly. This enables the identification of optimal treatment strategies based on true eradication events and allows us to explore how extinction probabilities depend on switching timing, antibiotic potency, and mutation rates.
@@
In [39], Beardmore and Peña-Miller state that a successful switching strategy, based on clinical observations [42], is “if the observed level of resistance to an antibiotic is too high, exchange it for a different antibiotic”, which is fully consistent with our results. In fact, a key result from our analysis is that fast sequential therapies are suboptimal, and that we need to allow for the evolution of resistance above a threshold in order to eradicate the population after the switch. The tension between waiting long enough for the single-resistant mutant to dominate the population and maximizing the number of switches leads to different optimal switching periods depending on the treatment duration (Fig. 6): longer therapies maximize extinction, but they also cause higher resistance, leading to a Pareto front of optimal switching periods. This gives us some additional insight that will help to choose optimal therapies.
@@
Furthermore, we show that, under our theoretical framework, we need strong CS for our therapies to work. While this seems a rather stringent condition, the good news is that such a therapy will work even under subinhibitory concentrations, which will always appear within the human body as a result of diffusion through tissues [43].
@@
In addition, we find a unimodal dose-extinction relationship (Fig. 5(c)), which has been observed before in both experiments [35] and models [44] and which can be fully explained using population dynamics arguments: we need single-resistant mutants to dominate the population; these are selected as antibiotic concentration increases, but selection can be hampered if doses are too high. We observe a similar relationship with mutation rates (Appendix F, Fig. 14), also supported by population dynamics arguments. This result suggests that increasing mutation rates might be a good complement to this kind of therapies, up to a point where we start facilitating the evolution of double-resistants. This suggestive strategy should be explored with care and checked experimentally before making any therapy recommendations.
This study adopts a deliberately minimal model, four genotypes and two antibiotics with symmetric effects, to isolate the mechanisms of CS-guided switching. Extensions to richer collateral-sensitivity networks [17, 18, 19] and to heterogeneous pharmacokinetics/pharmacodynamics [41] fit naturally within the same framework. CS relationships may vary during treatment [45, 46, 47]; such time-dependence can be incorporated by allowing parameters to evolve between switches and calibrating them from data. Empirical evaluation can proceed in vitro without altering the theoretical structure. A suitable test system is the ciprofloxacin–tobramycin pair in P. aeruginosa, where CS has been observed robustly and double resistance is infrequent [19, 48, 49], enabling direct assessment of the model’s predictions.
Beyond periodic schedules, our framework can be extended to nonperiodic and adaptive switching policies as considered in [39, 40]. Because we treat therapy as a series of switches, each a fresh chance to clear the infection, the same framework lends itself to control-theoretic optimization of switching sequences, including real-time, patient-specific adjustments informed by bacterial-load measurements. Overall, these results provide quantitative design rules for sequential therapy guided by collateral sensitivity and mathematical modeling, complementing current strategies against antimicrobial resistance.
Figure 8: Extinction histograms for the hybrid method: Tau-leaping and Gillespie’s algorithm. The top row (blue) shows extinction histograms with a total time of $t_{\text{end treatment}}=1000$ and $t_{\text{end}}=1050$ temporal units. As the antibiotic switching interval increases, extinction events become more clearly clustered after switching events (dashed black vertical lines). Dashed red vertical lines indicate $t_{\text{end treatment}}$. The middle row (pink) presents extinction histograms for a shorter total simulation time of $t_{\text{end treatment}}=100$ and $t_{\text{end}}=150$ temporal units, the same temporal scale used in the main text. At this scale, extinction events appear more diffusely distributed, making it difficult to discern their relationship with antibiotic changes. Both (blue and pink) consider extinction when the number of cells in the population is zero, and switch between tau-leaping and Gillespie’s algorithm with a threshold of 10 cells. The bottom row (green) shows results from Gillespie’s algorithm with a threshold of $0.05/\gamma$ for extinction and $15$ cells for switching between tau-leaping and Gillespie, the total simulation time of $t_{\text{end treatment}}=100$ and $t_{\text{end}}=150$ temporal units. The results are consistent to those observed with tau-leaping, supporting the strategy used throughout the main text.
Acknowledgments
@@
Software availability
The codes for our simulations, analyses, and production of figures are openly available [50].
Appendix A Justification for the population threshold and tau-leaping approximation
In the main text, we considered a stochastic model of bacterial population dynamics under antibiotic treatment, where extinction was defined as the population dropping below a threshold of $0.05\gamma^{-1}=5$ individuals. Simulations were performed using the tau-leaping method to efficiently approximate the stochastic process. Here, we justify this approach by comparing it with a hybrid method that uses tau-leaping for populations with many individuals and Gillespie’s algorithm for populations below that limit.
Figure 9: Changing the extinction threshold does not change the qualitative behavior of the extinction curve. Different thresholds for determining when a population is considered to be extinct, from 3% to 8% of theoretical carrying capacity. In the main text 5% has been used as the threshold.
Figure 10: (a) Cumulative extinction probability over 10 treatment switches with different switching periods $\tau$. Points represent extinction probabilities estimated through simulation, while the solid line reflects a fit to a geometric distribution $1-(1-p)^{n}$. (b) Estimation of the extinction probability for different number of antibiotic switches (represented with colors as indicated in the legend) using the fit of the geometric model obtained in (a). (c) Linear fit to the extinction probability after each switch for the hierarchical geometric model. (d) Estimated parameters for each tau $\tau$ in (c).
Gillespie’s algorithm provides an exact simulation of the underlying stochastic process but becomes computationally prohibitive when population size is large. Specifically, as the number of bacteria increases, the waiting times between reaction events shorten, leading to a significant increase in computational cost. This has a direct impact on the feasibility of simulating extinction events under different antibiotic switching strategies.
To assess whether the chosen threshold and tau-leaping approximation capture the relevant extinction dynamics, we performed simulations using the hybrid method and compared the results. Our key observation is that extinction events occur primarily after an antibiotic switch, with a characteristic delay. However, when antibiotic switching times are of the same order of magnitude as reaction times at low population sizes, this pattern is no longer evident. Instead, we observe a diffuse cloud of extinction events, making it difficult to distinguish the effect of antibiotic changes. To recover the expected correlation between switching events and extinctions, it is necessary to increase the switching times.
To illustrate these findings, we present extinction histograms from three sets of simulations. The first set (Fig. 8, top row) corresponds to a longer total simulation time of 1000 temporal units, allowing for a clearer observation of extinction clustering after antibiotic switches. The second set (Fig. 8, middle row) corresponds to a shorter total simulation time of 100 temporal units, matching the temporal scale used in the main text. In this shorter time frame, extinction events appear as a diffuse pattern, making it harder to discern correlations with antibiotic changes. In contrast, with a sufficiently long observation window, the extinction events align with antibiotic switching events, confirming that the primary driver of extinction is the switching strategy itself. In these two sets, we have considered extinctions when we have a population of zero bacteria. We add a third set using the hybrid method but with the same threshold for the extinctions used in the main text (Fig. 8, bottom row), confirming that the results are consistent and that the tau-leaping algorithm does not qualitatively alter the observed patterns.
We also tested the robustness of the extinction threshold. In the main text, we consider a population extinct if it goes below 5% carrying capacity. We varied this threshold between 3% and 8% and obtained quantitatively very similar results (Fig. 9).
Appendix B Parameters used in the main text
Unless otherwise stated, all graphs in the main text were made with the parameter values in Table 1.
Table 1: Model parameters and their meanings
Parameter Description Value
$\tau$ Duration of treatment with antibiotic $A$ or $B$ Variable
$k_{A}$, $k_{B}$ Effect of bacteriostatic antibiotics (higher $\Rightarrow$ weaker inhibition) 0.5
$\gamma$ Inverse of the theoretical carrying capacity 0.01
$k_{\rm CS}$ Collateral sensitivity effect (higher $\Rightarrow$ weaker CS) 0.05
$\mu_{1}$ Mutation rate to a resistant genotype $10^{-3}$
$\mu_{2}$ Mutation rate back to a sensitive genotype $10^{-2}$
$t_{\text{end}}$ Total time of the simulation 150
$t_{\text{end treatment}}$ Time in which we allow antibiotics to be switched 100
Appendix C Non-equilibrium effects and a hierarchical geometric model
Figure 11: Mean population trajectories. Mean values of the trajectories before the switch simulated for the geometric fit. Shades indicate 95% confidence interval. We observe that surviving populations achieve double-resistant mutants. (a) $N_{0}$. (b) $N_{3}$. (c) $N_{1}$ oscillatory behavior. (d) $N_{2}$ oscillatory behavior.
The discrepancies observed in the geometric model fit from Fig. 10(a,b) can be explained by the fact that the system is not in equilibrium. To better understand this behavior, we analyzed the extinction probability per round, $p$, as a function of time and the number of rounds. We performed a linear fit of the form $p=\alpha\tau+\beta n$, where $p$ is the extinction probability, $\tau$ is the switching period, and $n$ is the number of completed treatment rounds (Fig. 10(c)). This model captures the systematic dependence of extinction dynamics on both time and the structure of the treatment.
We then used this time-dependent extinction probability to construct a hierarchical geometric model for the cumulative extinction probability:
$P_{\text{ext}}=1-(1-(\alpha\tau+\beta n))^{n},$
which fits the data well over a wide range of conditions (Fig. 2(d)).
To explore how the fitting parameters vary with the switching period, we examined the dependence of $\alpha$ and $\beta$ on $\tau$ (Fig. 10(d)). For small values of $\tau$, we find $\beta>0$, indicating that the probability of extinction increases with the number of rounds. This could be due to the oscillatory regime driving the system into extinction-prone states. In contrast, for large $\tau$, we observe $\beta<0$, suggesting that prolonged exposure in each cycle may favor the emergence of double-resistant mutants, reducing extinction probability over time (Fig. 11). Together, these results highlight the importance of non-equilibrium effects in shaping extinction outcomes during sequential therapy.
Appendix D Parameters of the sigmoid fit
We introduce as parameters to be adjusted in a logistic model the population before the change of antibiotic. Depending on the antibiotic we are using, the populations of $x_{1}$ and $x_{2}$ have different growth rates, and therefore we consider the variables
$\begin{cases}x_{iA}=0\quad\text{if antibiotic is $B$,}\quad x_{i}\quad\text{otherwise}\\ x_{iB}=x_{i}\quad\text{if antibiotic is $B$,}\quad 0\quad\text{otherwise}\end{cases}$
for $i=1,2$. We therefore end up with the following inputs: $x_{0}$, $x_{3}$, $x_{1A}$, $x_{1B}$, $x_{2A}$, $x_{2B}$ (Table 2).
We observed that having $x_{0}$ or $x_{3}$ in the population before the switch reduces the probability of extinction (the parameters are negative). The same happens when we have bacteria that will be resistant after switching, $x_{1B}$ and $x_{2A}$. The best scenario for extinction is when all the population is dominated by either $x_{1A}$ or $x_{2B}$. The dataset used for these fits is the population composition before the switch of $10,000$ trajectories per $\tau$ simulated to estimate the extinction probability of Fig. 2(a).
Table 2: Parameters in sigmoid fit of Fig. 3
Parameter Sigmoid fit Sigmoid fit $\tau\geq 30$
$x_{0}$ -0.124302 -0.231164
$x_{3}$ -0.978612 -1.095430
$x_{1A}$ 0.014996 -0.004949
$x_{1B}$ -0.078326 -0.429122
$x_{2A}$ -0.035248 -0.512507
$x_{2B}$ 0.009665 -0.005999
Figure 12: Collateral sensitivity is irrelevant when high doses of antibiotic are applied. Simulations with $k=0.1$.
Figure 13: Effect of changes in antibiotic concentration. Mean population composition prior to the first switch (colored lines), superimposed with the extinction probability (black dashed line). As antibiotic inhibition increases ($k$ decreases), the population of $x_{1}$ goes up, which leads to an increase in extinction probability. After a certain antibiotic concentration (below $k\approx 0.7$), $x_{0}$ does not reach high values, and evolution to $x_{1}$ slows down. This leads to a decrease in extinction at intermediate to high antibiotic concentrations ($k<0.4$ approximately). When antibiotic inhibition becomes very strong (i.e. $k<0.1$), the decrease in $x_{0}$ is so strong that populations become extinct not due to collateral sensitivity, but because the overall carrying capacity is very low and extinctions occur from fluctuations. Simulations with $\tau=50$, $t_{\text{end}}=150$, $t_{\text{end treatment}}=100$, $\text{number of trajectories}=10000$.
Figure 14: Effect of changes in mutation rates. The extinction probability exhibits a nonmonotonic dependence on the baseline mutation rate ($\mu_{1}$), peaking at intermediate values. The ratio between mutation rates ($\mu_{2}/\mu_{1}$, color-coded) modulates the peak height and position. Simulations with $\tau=50$, $t_{\text{end}}=150$, $t_{\text{end treatment}}=100$, $\text{number of trajectories}=1000$.
Figure 15: Effect of changes in mutation rates. Mean population composition prior to the switch for trajectories that do not go extinct (colored lines), and extinction probability (black dashed line). The decline of the extinction peak coincides with the emergence of the double-resistant strain and with the recovery of the susceptible strain, the latter becoming more pronounced as $\mu_{2}$ increases relative to $\mu_{1}$. $\tau=50$, $t_{\text{end}}=150$, $t_{\text{end treatment}}=100$, $\text{number of trajectories}=1000$.
Appendix E Formula for the extension of the heuristic
Due to the boundary and initial conditions, Eq. (1) in the main text cannot be introduced directly into the geometric distribution for extending the extinction probability to multiple switches. Remember that $p(\tau)$ has two contributions, $p_{r}(\tau)$, the probability that the single-resistant mutant is not completely dominant in the population, and $p_{d}(\tau)$, the probability that a population gets extinct in a time smaller than $\tau$.
For the first antibiotic switch $p_{\mathrm{initial}}(\tau)$, we calculate $p_{r,\mathrm{initial}}(\tau)$ as
$p_{r,\mathrm{initial}}(\tau)=\text{Pr}(x_{0}+x_{2}+x_{3}\geq 2|\tau,x_{0}(0)=50)$
Similarly, after the last antibiotic switch, $p_{\mathrm{boundary}}(\tau)$, we modify the decay probability to $p_{d,\mathrm{boundary}}=p_{d}(t_{end}\text{ (mod }\tau+50))$, because the population has more time to decay, as a consequence of the chosen boundary conditions.
Putting everything together, we have the formula:
$p_{\mathrm{ext}}(\tau)=1-(1-p(\tau))^{n-2}\cdot(1-p_{\mathrm{initial}}(\tau))\cdot(1-p_{\mathrm{boundary}}(\tau))$
Appendix F Supplementary figures
Figs. 1215 contain supplementary information referred from the main text.